# FLoRA: Low-Rank Core Space for N-dimension

### v1

@@ 5.2 Baselines
We compare FLoRA with several state-of-the-art methods: fully fine-tuning, BitFit Zaken et al. (2021), HAdapter Houlsby et al. (2019), PAdapter Pfeiffer et al. (2020), AdaLoRA Zhang et al. (2022), the most representative low-rank method, LoRA Hu et al. (2021), and the state-of-the-art low-rank adaption method, DoRA Liu et al. (2024b). Specifically, HAdapter is strategically placed between the self-attention module and the FFN module, and it includes a subsequent residual connection. Conversely, PAdapter introduces a more streamlined design, implementing adapters exclusively after the FFN modules and LayerNorm modules. Furthermore, following Zhang et al. (2022), we apply AdaLoRA, LoRA and DoRA to all weight matrices or tensors. For more details on the baselines, please refer to their original papers.
@@ 5.3 Implementation Details
We compare FLoRA with other PEFT methods under different parameter budgets. The hidden dimension of Adapters is chosen from {8, 16, 32, 64}, the budget of AdaLoRA is chosen from {144, 288,567} and the rank $r$ of LoRA and DoRA is selected from {2, 4, 8, 16, 32}. Other hyper-parameters are initialized according to their original papers. Furthermore, we simply set $r_{1}=r_{2}$ for FLoRA. The scale $s$ is chosen from {0.04, 0.4, 4} for different foundation models. We use publicly available PyTorch Paszke et al. (2019) implementation to execute all the baseline comparisons, and all the experiments are conducted on NVIDIA A100 GPUs. For NLP tasks, the mean of 5 runs using various random seeds are reported for all the experiments, and all gains have passed the pairwise $t$-test with a significance of 0.05. The tensor $\mathcal{G}$ is initialized as zero, while other weight matrices are initialized as random Gaussian. For more training details, please refer to the Appendix.
@@
HAdapter 1.22M 90.13 95.53 68.64 91.27 94.11 84.48 89.95 91.48 88.19
PAdapter 1.18M 90.33 95.61 68.77 91.40 94.29 85.20 89.46 91.54 88.32
AdaLoRA 1.33M 90.38 95.87 71.45 91.19 94.36 88.09 90.69 91.84 89.23
LoRA 0.33M 90.03 93.92 69.15 90.61 93.37 85.56 90.19 90.75 87.95
DoRA 0.41M 90.21 94.38 69.33 90.84 93.26 86.94 90.19 91.34 88.31
FLoRA 0.33M 90.60 96.00 70.20 91.40 94.46 88.09 90.93 91.96 89.21
LoRA 1.33M 89.80 93.69 69.30 91.78 92.97 85.70 90.68 91.62 88.17
DoRA 1.41M 89.67 94.61 69.08 91.80 93.23 87.33 90.68 91.73 88.49
FLoRA 1.33M 90.82 96.21 72.05 91.94 94.60 89.53 91.18 92.04 89.80
@@ 5.4 Main Results
We compare FLoRA with other baselines under different parameter budgets. Specifically, ConvNeXt-V2-L is based on convolutions, and other large foundation models are dominantly based on linear layers. We evaluate FLoRA’s effectiveness on high-dimension space for the CV tasks, and on linear parameter space for CV and NLP tasks. The results are shown in Tables. 1-3 and Table in Appendix, and the normalized performance are illustrated in Fig. 2.
@@
For NLP tasks, FLoRA achieves better or on par performance compared with existing approaches on all datasets under all different parameter budgets. Specifically, under extreme low parameter budget, FLoRA performs better even than the baselines with higher parameter counts. For example, with 0.3M parameters, FLoRA’s performances on SST-2, QNLI, RTE, MRPC and STS-B are all better than baselines with larger parameter budgets.
@@ 6.1 Is the Core Space Truly Low-rank?
We present the performance of FLoRA under different rank (i.e., parameter budgets), tested on ConvNeXt-V2-L and DeBERTaV3-base. The results are illustrated in Fig. 3. It is evident that FLoRA’s performance with a low rank is comparable to, or even exceeds, that of a higher rank. This finding aligns with observations in Hu et al. (2021). It indicates that there indeed exists a core space in different dimensional spaces, and the rank of the core space is relatively small. When the rank set is smaller than the rank of the core space, the performance of the model is not optimal. Conversely, when it exceeds this rank, the core space is completely covered, which introduces some meaningless redundancy and noise. Additionally, we observe that for the convolutional parameter space, the rank of its core space is much larger, for the convolutional space has a more complex topological structure, necessitating a larger rank to adequately describe it.
@@ 6.2 Why is FLoRA’s Low-rank Representation Better than Other Methods?
Although methods like FLoRA and LoRA can represent changes in the original parameter space using low-rank representations, FLoRA performs better, indicating that its low-rank representation is superior to other forms of low-rank representations. We will explain the reasons as follows.
6.2.1 Theoretical Analysis
In linear parameter spaces, the changes modeled by methods in the LoRA derivatives, can also be simply considered as $\mathbf{AGB}$ with $\mathbf{G}$ being a diagonal matrix. However, FLoRA does not impose any constraints on $\mathbf{G}$, which offers several advantages: firstly, it broadens the range of parameter adjustments and enhances the flexibility of parameter learning. Secondly, FLoRA removes the strong orthogonality constraint on $\mathbf{G}$. Since orthogonality is not a characteristic of all downstream tasks, enforcing orthogonal constraints during the learning process could potentially degrade performance, as shown in Tables 1-3. Lastly, since $\mathbf{G}$ is initialized to zero in FLoRA, under certain conditions, $\mathbf{G}$ may also learn to exhibit orthogonal properties. This indicates that the low-rank representations of methods like LoRA are a special case of FLoRA, which possesses a more robust low-rank representation capability, thereby yielding better performance.
@@ 6.2.1 Theoretical Analysis
6.2.2 Empirical Analysis
To further substantiate the theoretical analysis on FLoRA’s superiority, we record the average of the Frobenius norm of $\Delta\mathbf{W}$ and the feature amplification factor $\frac{\|\Delta\mathbf{W}\|_{F}}{\|\mathbf{U}^{\mathsf{T}}\mathbf{W}\mathbf{V}^{\mathsf{T}}\|_{F}}$ Hu et al. (2021) of all the layers in DeBERTaV3-base on CoLA during training. Here $\mathbf{U}$ and $\mathbf{V}$ being the left- and right-singular matrices of the SVD decomposition of $\Delta\mathbf{W}$. The feature amplification factor measures how much of task-specific information are amplified by $\Delta\mathbf{W}$.
The results, illustrated in Fig. 4, show that LoRA and DoRA can amplify task-specific features more than FLoRA in the early stages. As discussed in Sec. 6.2.1, these two methods exhibit strong orthogonality, resulting in a distinct directional learning pattern at the beginning, which contributes to their larger initial values. However, strong orthogonality may not always be suitable for downstream tasks, and since downstream datasets may possess various properties, their amplification factors upon convergence are smaller than that of FLoRA.
Additionally, we found the trend of the Frobenius norm closely aligns with the feature amplification factor, surprisingly. This might indicate that larger $\Delta\mathbf{W}$ values can accommodate more task-specific information, thereby more effectively amplifying the task-specific information in the frozen weights. We also observed that, initially, the Frobenius norms of $\Delta\mathbf{W}$ for LoRA and DoRA are also larger than that of FLoRA, for they can quickly capture the orthogonal information if the downstream datasets. However, upon convergence, they are both smaller than FLoRA’s, suggesting that FLoRA can contain more kinds of properties for task-specific information.
@@ 6.4 Is FLoRA Sensitive to the Scaling Factor?
7 Conclusion and Limitation
@@
In this paper, we propose a generalized low-rank tensor decomposition based PEFT method, FLoRA, aiming for N-dimension parameter space. FLoRA asserts that the alternations in each dimensional parameter space contain a low-rank core space structurally consistent to the original space. It models updates using this core space alongside corresponding weights to reconstruct the original alternation space. By doing so, FLoRA effectively preserves the structural integrity of the original N-dimensional parameter space while decomposing it through low-rank tensor decomposition. Extensive experiments on computer vision, natural language processing and multi-modal domains substantiate the effectiveness of FLoRA.

### v2

@@ 4.2 Baselines
We compare FLoRA with several state-of-the-art methods: fully fine-tuning, BitFit Zaken et al. (2021), HAdapter Houlsby et al. (2019), PAdapter Pfeiffer et al. (2020), AdaLoRA Zhang et al. (2022a), the most representative low-rank method, LoRA Hu et al. (2021a), and the state-of-the-art low-rank adaption method, DoRA Liu et al. (2024b). Specifically, HAdapter is strategically placed between the self-attention module and the FFN module, and it includes a subsequent residual connection. Conversely, PAdapter introduces a more streamlined design, implementing adapters exclusively after the FFN modules and LayerNorm modules. Furthermore, following Zhang et al. (2022a), we apply AdaLoRA, LoRA and DoRA to all weight matrices or tensors. For more details on the baselines, please refer to their original papers.
@@ 4.3 Implementation Details
We compare FLoRA with other PEFT methods under different parameter budgets. The hidden dimension of Adapters is chosen from {32, 64}, the budget of AdaLoRA is set as {144, 288, 576} and the rank $r$ of LoRA and DoRA is selected from {2, 4, 8, 16, 32}. Other hyper-parameters are initialized according to their original papers.
Additionally, we simply set $r=r_{1}=r_{2}$ for FLoRA. For CV tasks, when fine-tuning ConvNeXt-V2-L, we set $r_{3}=2$ and $s=4$. We adjusted FLoRA’s $r$ to align its parameter budget closely with that of other methods for fair comparison. For different methods, we fine-tune all the convolutional layers for ConvNeXt-V2-L (and SDXL). When fine-tuning InternViT-6B, FLoRA’s $r$ is set to {16, 32}, $s=0.04$, and we fine-tune all the linear layers for different methods. For NLP tasks, FLoRA’s $r$ is set to {2, 8}, and $s=0.4$. The mean of 5 runs using various random seeds are reported for all the experiments, and all gains have passed the pairwise $t$-test with a significance of 0.05. We fine-tune q, k, v, up and down matrices for all methods. For MM tasks, we set $r=128$, $s=1.5$ and fine-tune all linear layers. More training details can be found in Appendix.
We use publicly available PyTorch Paszke et al. (2019) implementation to execute all the baseline comparisons, and all the experiments are conducted on NVIDIA A100 GPUs. The tensor $\mathcal{G}$ is initialized as zero, while other weight matrices are initialized as random Gaussian. For more training details, please refer to the Appendix.
@@
HAdapter 1.22M 90.13±.2 95.53±.1 68.64±.4 91.27±.0 94.11±.1 84.48±.9 89.95±1.1 91.48±.3 88.19
PAdapter 1.18M 90.33±.3 95.61±.3 68.77±.9 91.40±.1 94.29±.2 85.20±1.2 89.46±.7 91.54±.7 88.32
AdaLoRA 1.33M 90.38±.3 95.87±.3 71.45±1.2 91.19±.2 94.36±.3 88.09±1.9 90.69±1.5 91.84±.7 89.23
LoRA 0.33M 90.03±.3 93.92±.2 69.15±1.2 90.61±.1 93.37±.3 85.56±.7 90.19±.7 90.75±.2 87.95
DoRA 0.41M 90.21±.2 94.38±.5 69.33±1.9 90.84±.1 93.26±.3 86.94±.7 90.19±1.2 91.34±.9 88.31
FLoRA 0.33M 90.60±.2 96.00±.5 70.20±1.1 91.40±.4 94.46±.6 88.09±.5 90.93±.6 91.96±.7 89.21
LoRA 1.33M 89.80±.1 93.69±.2 69.30±1.1 91.78±.1 92.97±.1 85.70±.4 90.68±.6 91.62±.3 88.17
DoRA 1.41M 89.67±.2 94.61±.5 69.08±1.3 91.80±.1 93.23±.2 87.33±1.1 90.68±1.0 91.73±.5 88.49
FLoRA 1.33M 90.82±.1 96.21±.3 72.05±.3 91.94±.1 94.60±.1 89.53±.2 91.18±.3 92.04±.3 89.80
Table 4: Results with LLaVA-1.5-7B Liu et al. (2024a) fine-tuned on visual instruction tuning.
Method Params (%) VQAv2 GQA VisWiz SQA VQAT POPE MMBench Avg.
Fully FT 100 78.5 61.9 50.0 66.8 58.2 85.9 64.3 66.5
LoRA 4.61 79.1 62.9 47.8 68.4 58.2 86.4 66.1 66.9
DoRA 4.63 78.6 62.9 52.2 69.9 57.0 87.2 66.1 67.6
FLoRA 4.63 78.6 62.9 51.7 70.1 57.4 87.6 66.0 67.8
@@ 4.4 Main Results
We compare FLoRA with other baselines under different parameter budgets. Specifically, ConvNeXt-V2-L is based on convolutions, and other large foundation models are dominantly based on linear layers. We evaluate FLoRA’s effectiveness on high-dimension space for the CV tasks, and on linear parameter space for CV and NLP tasks. The results are shown in Tables. 1-4 and Tables. 7-8 in Appendix, and the normalized performance are illustrated in Fig. 2.
@@
For NLP tasks, FLoRA achieves better or on par performance compared with existing approaches on all datasets under all different parameter budgets. Specifically, under extremely low parameter budget, FLoRA performs better even than the baselines with higher parameter counts. For example, with 0.3M parameters, FLoRA’s performances on SST-2, QNLI, RTE, MRPC and STS-B are all better than baselines with larger parameter budgets.
@@ 5.1 Is the Core Space Truly Low-rank?
We present the performance of FLoRA under different rank (i.e., parameter budgets), tested on ConvNeXt-V2-L and DeBERTaV3-base. The results are illustrated in Fig. 3. It is evident that FLoRA’s performance with a low rank is comparable to, or even exceeds, that of a higher rank. This finding aligns with observations in Hu et al. (2021a). It indicates that there indeed exists a core space in different dimensional spaces, and the rank of the core space is relatively small. When the rank set is smaller than the rank of the core space, the performance of the model is not optimal. Conversely, when it exceeds this rank, the core space is completely covered, which introduces some meaningless redundancy and noise. Additionally, we observe that for the convolutional parameter space, the rank of its core space is much larger, for the convolutional space has a more complex topological structure, necessitating a larger rank to adequately describe it.
@@ 5.2 Why is FLoRA’s Low-rank Representation Better than Other Methods?
Although methods like FLoRA and LoRA can represent changes in the original parameter space using low-rank representations, FLoRA performs better, indicating that its low-rank representation is superior to other forms of low-rank representations. We will explain the reasons as follows. We also provide a comprehensive analysis in Appendix B.1.
In linear parameter spaces, the changes modeled by methods in the LoRA derivatives, can also be simply considered as $\mathbf{AGB}$ with $\mathbf{G}$ being a diagonal matrix. However, FLoRA does not impose any constraints on $\mathbf{G}$, i.e., matrix pattern, which offers several advantages: firstly, it broadens the range of parameter adjustments and enhances the flexibility of parameter learning. Secondly, FLoRA removes the constraint on $\mathbf{G}$. Since any matrix patterns may not represent the characteristics of all downstream tasks, enforcing these constraints on matrix patterns during the learning process could potentially degrade performance. as shown in Tables 1-3. Moreover, Si et al. (2024b) has also theoretically validates the effectiveness of FLoRA, which is significantly easier to reach the global optimum compared to LoRA. This indicates that the low-rank representations of methods like LoRA are inferior to that of FLoRA, thereby yielding inferior performances.
@@
To further substantiate the analysis on FLoRA’s superiority, we record the average of the Frobenius norm of $\Delta\mathbf{W}$ and the feature amplification factor $\frac{\|\Delta\mathbf{W}\|_{F}}{\|\mathbf{U}^{\mathsf{T}}\mathbf{W}\mathbf{V}\|_{F}}$ Hu et al. (2021a) of all the layers in DeBERTaV3-base on CoLA during training, with FLoRA, DoRA and LoRA’s rank $r=8$. Here $\mathbf{U}$ and $\mathbf{V}$ being the top $r$ left- and right-singular matrices of the SVD decomposition of $\Delta\mathbf{W}$. The feature amplification factor measures how much of task-specific information are amplified by $\Delta\mathbf{W}$.
The results, illustrated in Fig. 4, show that LoRA and DoRA can amplify more task-specific features than FLoRA in the early stages. As discussed before, these two methods have strong constraints on matrix patterns, resulting in a distinct directional learning pattern at the beginning, which contributes to their larger initial values. However, such constraints on matrix patterns may not always be suitable for downstream tasks, and since downstream datasets may possess various properties, their amplification factors upon convergence are smaller than that of FLoRA.
Additionally, we found the trend of the Frobenius norm closely aligns with the feature amplification factor, surprisingly. This might indicate that a larger Frobenius norm of $\Delta\mathbf{W}$ values can accommodate more task-specific information, thereby more effectively amplifying the task-specific information in the frozen weights. We also observed that, initially, the Frobenius norms of $\Delta\mathbf{W}$ for LoRA and DoRA are also larger than that of FLoRA, for they can quickly capture the information due to the learning of specific matrix patterns. However, upon convergence, they are both smaller than FLoRA’s, suggesting that FLoRA can contain more kinds of properties for task-specific information.
@@ 5.4 Is FLoRA Sensitive to the Scaling Factor?
6 Related Work
6.1 Parameter-Efficient Fine-tuning
Methods for Parameter-Efficient Fine-tuning (PEFT) have been conceived to mitigate the substantial computational costs associated with the fine-tuning of large-scale models. This economization is realized by honing a comparatively minute fraction of the overall parameters, selected strategically for adaptation to a variety of downstream tasks. Current PEFT techniques can be divided into three distinct categories Liu et al. (2024b); Ding et al. (2023): Adapter-based Houlsby et al. (2019); Chen et al. (2022); Luo et al. (2023); He et al. (2021a); Mahabadi et al. (2021); Karimi Mahabadi et al. (2021), prompt-based Lester et al. (2021); Razdaibiedina et al. (2023); Wang et al. (2023b); Shi & Lipani (2023); Fischer et al. (2024) and low-rank matrix adaptation-based Hu et al. (2021a); Liu et al. (2024b); Hyeon-Woo et al. (2021); Qiu et al. (2023); Renduchintala et al. (2023); Kopiczko et al. (2023); YEH et al. (2023); Zhang et al. (2022a); Si et al. (2024a). The first category of method integrates linear modules either sequentially or concurrently with the existing layer to enhance the performance, and the second class introduces additional soft tokens (prompts) to the initial input and concentrate exclusively on refining these trainable vectors. The last type, proposed by LoRA Hu et al. (2021a), adopts low-rank matrix adaptation to model the change of the weight during fine-tuning, and are capable of merging with pre-trained weights.
However, these methods only focuses on linear weights or destructing the structural integrity of high dimensional parameter spaces. To this end, we propose a novel method FLoRA to address various dimensional parameter space.
6.2 Tucker Decomposition
When designing FLoRA, our goal was to preserve the topological structure of different dimensions of parameter space while maintaining the changes within a low-rank manifold. To achieve this, we use a tensor of the same dimensionality as the original parameter space to model its changes. By learning the corresponding projection matrices, we align the dimensional sizes of the low-rank space with those of the original space. From this perspective, we found that the form of Tucker decomposition Tucker (1966) naturally aligns with our approach.
Tucker decomposition is one of the well-studied algebraic tensor decomposition. Formally, given a tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$, where $N$ is the order of the tensor (i.e., the number of dimensions or modes), Tucker decomposition represents $\mathcal{X}$ as a product of a core tensor $\mathcal{G}\in\mathbb{R}^{J_{1}\times J_{2}\times\cdots\times J_{N}}$ and a matrix along each mode $n$, $\mathbf{A}^{(n)}\in\mathbb{R}^{I_{n}\times J_{n}}$, where $J_{n}$ can be considered as the dimension of the core tensor along mode $n$. The decomposition can be compactly written as:
where ×$\times_{n}$ denotes the or mode-$n$ product between a tensor and a matrix. The core tensor $\mathcal{G}$ represents the interactions between different modes, and the matrices $\mathbf{A}^{(n)}$ are analogues to principal components within each respective mode. The selection of dimensions $J_{1},J_{2},\ldots,J_{N}$ allows for a balance between desired approximation quality and computational efficiency, tailored to the specific requirements of the task at hand.
6.3 Other Methods Based on Low-rank Tensor Adaptations
There are some works like LoTR Bershatsky et al. (2024) and SuperLoRA Chen et al. (2024b) also adopt low-rank tensor adaptations, but they are unrelated to FLoRA.
Usage and Impact. Based on low-rank tensor adaptation, FLoRA directly models $\Delta\mathbf{W}$ for each weight, maintaining its spatial structure. However, LoTR concatenates $\Delta\mathbf{W}$ across all layers and then applies low-rank tensor adaptation, assuming all layers share same weight matrices, which is a strong assumption that leads to suboptimal performance. SuperLoRA concatenates the vectorized $\Delta\mathbf{W}$ across all Multi-Layer Attention layers, performs various operations, and then redistributes them back to the original dimensions. In this process, low-rank tensor adaptation is just one choice. Besides, both LoTR and SuperLoRA disrupt the original parameter space structure although they use low-rank tensor adaptation.
Motivation and Details. FLoRA aims to model $\Delta\mathbf{W}$ while preserving the N-dimensional topological structure, meanwhile addressing the limitations of other methods focusing only on linear weights. However, LoTR aims to achieve extreme parameter efficiency, and SuperLoRA aims to unify and extend LoRA derivatives by setting different hyper-parameters. Both LoTR and SuperLoRA are tailored for linear weights, limiting their applications.
7 Conclusion
@@
In this paper, we propose a generalized low-rank tensor adaptation based PEFT method, FLoRA, aiming for N-dimension parameter space. FLoRA asserts that the alternations in each dimensional parameter space contain a low-rank core space structurally consistent to the original space. It models updates using this core space alongside corresponding weights to reconstruct the original alternation space. By doing so, FLoRA effectively preserves the structural integrity of the original N-dimensional parameter space while modeling its change through low-rank tensor adaptations. Extensive experiments on computer vision, natural language processing and multi-modal domains substantiate the effectiveness of FLoRA.